An electric dipole consist of point charges q1=+12nC and q2= -12nC placed 10.0 cm apart. Compute the electric potential energy.


1
Expert's answer
2022-03-02T10:49:07-0500

The electrical potential energy

"W=k\frac{q_1q_2}{r}\\
=9*10^9*\frac{12*10^{-9}*(-12*10^{-9})}{0.1}=-1.3*10^{-5}\:\rm J"